OpenStudy (anonymous):

(-9)÷(1-((-4)-(-8)))

OpenStudy (anonymous):

-9/(1-4) -9/-3=3

OpenStudy (anonymous):

whats the steps to solve it

OpenStudy (anonymous):

go to wolfram alpha. it will give you all the steps

OpenStudy (anonymous):

start on the inside -4--8=-4+8=4 1-4=-3 -9/-3=3

OpenStudy (anonymous):

(-9) / (1- (-4+8)) = (-9) / (1-(4)) = (-9) / (-3) = 3